OpenStudy (anonymous):

Whats is the Interval Notation of x ≥1 or ≥-3

OpenStudy (jdoe0001):

http://www.mathsisfun.com/sets/intervals.html \(\bf x \ge 1\\ x \ge -3\) notice you're using \(\bf \ge \) that means it will include the endpoint, so it'll be a solid dot there

OpenStudy (anonymous):

This is what I got [1,infinite)U(Infinite,-3}

OpenStudy (jdoe0001):

you mean \(\bf \large (-\infty\quad , \quad q\quad -3) \cup (1\quad , \quad +\infty) \quad ?\)

OpenStudy (anonymous):

{1,∞) U (∞,-3}

OpenStudy (jdoe0001):

woops, I have an extra character there

OpenStudy (jdoe0001):

hmmm \(\bf \large (-\infty\quad , \quad \quad -3) \cup (1\quad , \quad +\infty) \quad ?\) rather

OpenStudy (jdoe0001):

well... ahemm... on the negative side of the line as you get closer to the 0, the value of the negative amount is BIGGER so -2 is really greater than -3 while 2 is lesser than 3 on the positive side so \(\bf [-3\quad , \quad +\infty) \cup [1\quad , \quad +\infty)\\ \textit{then again, the 1st interval will cover the 2nd one}\\ \textit{so the 1st interval will suffice, that is }\large [-3\quad , \quad +\infty) \)

OpenStudy (anonymous):

I think the answer shuld be {1,∞)U(∞,-3)

OpenStudy (jdoe0001):

I see the infinity symbol there, but there are 2 infinities, + and - do you mean the positive one?

OpenStudy (anonymous):

Oh I got it

OpenStudy (anonymous):

What about x ≥2 or x ≥-3 help me draw the number line and internal notation

OpenStudy (jdoe0001):

anything that's greater than or equal to -3, is also greater than or equal to +2 so one can say \(\bf x \ge 2\qquad x \ge -3\\ \quad \\ [-3\quad , \quad +\infty)\cup [2\quad , \quad +\infty) \implies [-3\quad , \quad +\infty)\)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

Thanks ..

OpenStudy (anonymous):

what about plotin this on a number line [−3,+∞)u[1,+∞)

OpenStudy (jdoe0001):

if you notice the interval for that one, is the same graph